The night of September 24 was the first night of LBTO’s 2013B observing semester.

After an intense shutdown which ended at the end of August, LBTO staff has been very busy checking out many telescope subsystems and instruments, which went through maintenance or upgrades, as well as the adaptive secondaries (especially AdSec-DX).

Thanks to a generous engineering time allocation devoted to these efforts, much was accomplished with a decent end of monsoon weather (much better than in 2011 and 202!) A good thing, as there was much to do this year before being able to go back to science!

The first night of 13B was a good night science wise with minimum glitches. A relief after the few problems encountered on instruments at the last minute, which brought some interesting moments (!) to those involved. The LBCs, LUCI1 and AO on both sides are operational. MODS1 has an issue on its red channel (likely a controller problem), which is being addressed by LBTO and the OSU team. Hopefully, MODS1 will be back soon.

The current version of the 2013B telescope schedule is available here

We will use this blog to give news on the status of the observatory as soon as things happen. The last posts show up on the LBTO web page.

The posts are broadcasted on the LBT Facebook page

They are also linked on Twitter
Follow @LBTObs

Finally, you can subscribe to the blog feed through the link given at the bottom of this page.